# The economics of global personality diversity

### v3

@@ Results
Figure 1: |||| Hierarchical clustering of Big Five personality facets across 135 countries. The dendrogram represents the clustering of countries based on median scores across 30 Big Five personality dimensions, using data from a global sample (n = 760,242). The analysis reveals significant variation in personality profiles across countries, with patterns linked to shared cultural, economic, historical, and geographic contexts. While familiar regional pairings are evident, the clustering also highlights unexpected and intriguing combinations, suggesting complex interplays between personality and national characteristics.
@@
Figure 2: |||| In-country personality combinations cluster into distinct functional types with economic significance. This dendrogram and heatmap illustrate eight personality clusters (Pathfinder, Leader, Advisor, Translator, Yearner, Pioneer, Strategist, Uniter) based on Big Five personality traits across 30 dimensions (n = 44,678), using the United States as an example. The clusters reveal the structure of the PLATYPUS model, showing clear differences in personality profiles associated with varying economic roles and capabilities. Entrepreneurial personalities are notably concentrated in the Pathfinder and Leader clusters, which are linked to founders of startups, innovative and and high-growth firms. Supplementary Information provides the distribution of predicted entrepreneurial personalities by cluster.
@@
Type Archetype Description Key Big Five Personality Facets Economic Functions Pathfinder Proactive Collaborators High in Agreeableness (Trust); Openness (Adventurousness) and Extraversion (Activity Level) Build cohesive teams and foster an inclusive, collaborative culture. Leader Charismatic Leaders High in Agreeableness (Co-ooperation and Morality) and Conscientiousness (Dutifulness). Foster collaboration and drive sustainable growth. Advisor Empathetic Guides High in Conscientiousness (Dutifulness) and Agreeableness (Morality and Altruism) Maintain social infrastructure, essential for productivity. Translator Optimistic Connectors High in Extraversion (Cheerfulness, Excitement-Seeking, Gregariousness) Creating cohesive social and professional networks. Yearner Emotional Idealists High in Openness (Imagination, Emotionality) and Emotional Stability (Anxiety) Offering unique insights into human experiences. Yearner Creative Empaths High in Openness (Artistic Interests) and Low Conscientousness (Cautiousness) Foster innovation and social well-being through creativitiy and spontaneity. Uniter Balanced Dependables Moderate in all Big Five traits, showing versatility and flexibility Bridge gaps between creativity and practical implementation. Strategist Practical Realists High in Conscientiousness (Orderliness, Cautiousness), Low in Agreeableness (Sympathy) Thriving in roles that demand precision and careful planning.
@@
Figure 3: |||| Global Personality Diversity Index ($\Psi$-GPDI) distributions across 135 countries. Ridgeline plot illustrating the Global Personality Diversity Index ($\Psi$-GPDI) for 135 countries, based on pairwise cosine similarity distances in personality traits. The $\Psi$-GPDI reveals the extent of personality diversity within each country, showcasing variations in median values, distribution shapes, and multimodal patterns. These findings highlight distinct national personality dynamics, with implications for cultural cohesion, innovation potential, and economic outcomes.
@@
Figure 4: |||| GDP per capita and Personality Diversity Index ($\Psi$-GPDI) across 135 countries. Scatter plot showing the relationship between the Personality Diversity Index ($\Psi$-GPDI) and GDP per capita (in constant 2005–2023 PPP \$) for 135 countries. A positive association is observed, with higher personality diversity correlating with greater economic productivity in most regions. However, deviations such as a plateau or decline in some cases highlight the complex interplay between diversity and economic outcomes. Labels identify select OECD and G20 countries. This analysis underscores the potential role of personality diversity in driving national income and innovation metrics.
@@
Figure 5: |||| Multifactor analysis identifying Personality Diversity Index ($\Psi$-GPDI) as a key third factor in explaining national income. Scatterplots and correlation matrix illustrating the relationships between GDP per capita, Personality Diversity Index ($\Psi$-GPDI), migration, and expropriation risk. The analysis positions $\Psi$-GPDI as a third key factor, alongside two established factors—migration (Corr = 0.488) and expropriation risk (Corr = 0.716)—in explaining variance in national income. The moderate correlation between $\Psi$-GPDI and GDP (Corr = 0.453) highlights its potential as a novel explanatory variable in understanding economic outcomes, complementing demographic and economic risk factors.
Table 2: |||| Contributions of Expropriation, Migration, and Personality to GDP, showing the adjusted R-squared values for individual and combined factors. The table also includes Variance Inflation Factor (VIF) values for each factor, assessing multicollinearity. The associated Venn diagram illustrates the overlapping contributions of these factors to national income, highlighting the links between personality diversity, migration, expropriation risk, and GDP.
@@ A.1 Data representativeness
Extended Data Fig. 1: |||| Comparison of Johnson Data Age Distribution and Global Population (UN 2010). The age distribution of the Johnson sample (blue bars) is compared to the global population distribution (red line) based on UN 2010 data. The Johnson sample exhibits a linear correlation (r = 0.64, p = 0.000000001%), supporting its representativeness. A logarithmic transformation is applied to the global population data to align scales and illustrate parallel trends, highlighting the consistency in relative age distribution patterns between the two datasets.
Extended Data Fig. 2: |||| Comparison of Johnson Data Country Distribution and Global Population by Country (UN 2010). The country distribution of the Johnson sample (light blue bars) is compared to the global population distribution by country (red line) based on UN 2010 data, with China and India omitted for clarity. The two datasets show a linear correlation (r = 0.54, p = 0.000007811%), indicating the representativeness of the Johnson sample. The logarithmic trend highlights the differences in scale while preserving the general alignment of patterns across countries.
@@ A.2 Equality test of pairwise cosine similarity distributions of 135 countries
Extended Data Fig. 3: |||| Global personality diversity distribution comparison. The Kolmogorov-Smirnov (KS) test reveals that most countries have significantly different distributions of personality diversity.
@@ A.3 Relationships between PDI and GDP
Extended Data Fig. 4: |||| A, OECD + G20 GDP per capita for “Magic 46”. B, GDP per capita only African Countries.

### v4

@@ Results
Figure 1: $|$ Hierarchical clustering of Big Five personality facets across 135 countries. The dendrogram represents the clustering of countries based on median scores across 30 Big Five personality dimensions, using data from a global sample (n = 760,242). The analysis reveals significant variation in personality profiles across countries, with patterns linked to shared cultural, economic, historical, and geographic contexts. While familiar regional pairings are evident, the clustering also highlights unexpected and intriguing combinations, suggesting complex interplays between personality and national characteristics.
@@
Figure 2: $|$ In-country personality combinations cluster into distinct functional types with economic significance. This dendrogram and heatmap illustrate eight personality clusters (Pathfinder, Leader, Advisor, Translator, Yearner, Pioneer, Strategist, Uniter) based on Big Five personality traits across 30 dimensions (n = 44,678), using the United States as an example. The clusters reveal the structure of the PLATYPUS model, showing clear differences in personality profiles associated with varying economic roles and capabilities. Entrepreneurial personalities are notably concentrated in the Pathfinder and Leader clusters, which are linked to founders of startups, innovative and and high-growth firms. Supplementary Information provides the distribution of predicted entrepreneurial personalities by cluster.
@@
Type Archetype Description Key Big Five Personality Facets Economic Functions
Pathfinder Proactive Collaborators High in Agreeableness (Trust); Openness (Adventurousness) and Extraversion (Activity Level) Build cohesive teams and foster an inclusive, collaborative culture.
Leader Charismatic Leaders High in Agreeableness (Co-ooperation and Morality) and Conscientiousness (Dutifulness). Foster collaboration and drive sustainable growth.
Advisor Empathetic Guides High in Conscientiousness (Dutifulness) and Agreeableness (Morality and Altruism) Maintain social infrastructure, essential for productivity.
Translator Optimistic Connectors High in Extraversion (Cheerfulness, Excitement-Seeking, Gregariousness) Creating cohesive social and professional networks.
Yearner Emotional Idealists High in Openness (Imagination, Emotionality) and Emotional Stability (Anxiety) Offering unique insights into human experiences.
Yearner Creative Empaths High in Openness (Artistic Interests) and Low Conscientousness (Cautiousness) Foster innovation and social well-being through creativitiy and spontaneity.
Uniter Balanced Dependables Moderate in all Big Five traits, showing versatility and flexibility Bridge gaps between creativity and practical implementation.
Strategist Practical Realists High in Conscientiousness (Orderliness, Cautiousness), Low in Agreeableness (Sympathy) Thriving in roles that demand precision and careful planning.
@@
Figure 3: $|$ Global Personality Diversity Index ($\Psi$-GPDI) distributions across 135 countries. Ridgeline plot illustrating the Global Personality Diversity Index ($\Psi$-GPDI) for 135 countries, based on pairwise cosine similarity distances in personality traits. The $\Psi$-GPDI reveals the extent of personality diversity within each country, showcasing variations in median values, distribution shapes, and multimodal patterns. These findings highlight distinct national personality dynamics, with implications for cultural cohesion, innovation potential, and economic outcomes.
@@
Figure 4: $|$ GDP per capita and Personality Diversity Index ($\Psi$-GPDI) across 135 countries. Scatter plot showing the relationship between the Personality Diversity Index ($\Psi$-GPDI) and GDP per capita (in constant 2005–2023 PPP \$) for 135 countries. A positive association is observed, with higher personality diversity correlating with greater economic productivity in most regions. However, deviations such as a plateau or decline in some cases highlight the complex interplay between diversity and economic outcomes. Labels identify select OECD and G20 countries. This analysis underscores the potential role of personality diversity in driving national income and innovation metrics.
@@
Figure 5: $|$ Multifactor analysis identifying Personality Diversity Index ($\Psi$-GPDI) as a key third factor in explaining national income. Scatterplots and correlation matrix illustrating the relationships between GDP per capita, Personality Diversity Index ($\Psi$-GPDI), migration, and expropriation risk. The analysis positions $\Psi$-GPDI as a third key factor, alongside two established factors—migration (Corr = 0.488) and expropriation risk (Corr = 0.716)—in explaining variance in national income. The moderate correlation between $\Psi$-GPDI and GDP (Corr = 0.453) highlights its potential as a novel explanatory variable in understanding economic outcomes, complementing demographic and economic risk factors.
Table 2: $|$ Contributions of Expropriation, Migration, and Personality to GDP, showing the adjusted R-squared values for individual and combined factors. The table also includes Variance Inflation Factor (VIF) values for each factor, assessing multicollinearity. The associated Venn diagram illustrates the overlapping contributions of these factors to national income, highlighting the links between personality diversity, migration, expropriation risk, and GDP.
@@ A.1 Data representativeness
Extended Data Fig. 1: $|$ Comparison of Johnson Data Age Distribution and Global Population (UN 2010). The age distribution of the Johnson sample (blue bars) is compared to the global population distribution (red line) based on UN 2010 data. The Johnson sample exhibits a linear correlation (r = 0.64, p = 0.000000001%), supporting its representativeness. A logarithmic transformation is applied to the global population data to align scales and illustrate parallel trends, highlighting the consistency in relative age distribution patterns between the two datasets.
Extended Data Fig. 2: $|$ Comparison of Johnson Data Country Distribution and Global Population by Country (UN 2010). The country distribution of the Johnson sample (light blue bars) is compared to the global population distribution by country (red line) based on UN 2010 data, with China and India omitted for clarity. The two datasets show a linear correlation (r = 0.54, p = 0.000007811%), indicating the representativeness of the Johnson sample. The logarithmic trend highlights the differences in scale while preserving the general alignment of patterns across countries.
@@ A.2 Equality test of pairwise cosine similarity distributions of 135 countries
Extended Data Fig. 3: $|$ Global personality diversity distribution comparison. The Kolmogorov-Smirnov (KS) test reveals that most countries have significantly different distributions of personality diversity.
@@ A.3 Relationships between PDI and GDP
Extended Data Fig. 4: $|$ A, OECD + G20 GDP per capita for “Magic 46”. B, GDP per capita only African Countries.